OpenStudy (anonymous):

solve for all real values of x. (x^2)-9x+9=0

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

\[x=\frac{9\pm3\sqrt{5}}{2}\]

OpenStudy (anonymous):

method is formula, not way easier. i will write it if you like

OpenStudy (sriram):

9-sqrt45 9+sqrt45 ------- or ------- 2 2

OpenStudy (anonymous):

x^2-9x+9=0 Okay, so we start by using the quadratic formula, \[(-b \pm \sqrt{b ^{2}-4ac})/2a\] This gives solutions as follows: \[9/2 \pm (3/2)\sqrt{5}\] Hope this helps. http:/www.tutorsean.net